# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 1 Introduction
This survey (re)introduces reinforcement learning to economists. I review the theoretical connections between dynamic programming and reinforcement learning, demonstrating how value iteration, Q-learning, and policy gradient methods are common solution methods to the same class of optimization problems. I then examine applications across several domains, including control problems including pricing and inventory management; structural economic models with high-dimensional state spaces; strategic games in which multi-agent algorithms compute equilibria under imperfect information; bandit problems in which economic structure yields tighter regret bounds; and preference learning. The exposition combines formal theory, practical applications and computational illustration.
Both dynamic programming and reinforcement learning solve the Bellman equation; they differ in the information requirements and the way in which the solution is refined. First, dynamic programming requires knowledge of the transition in the environment and the reward function which allows the reduction of the average Bellman error, reinforcement learning estimates value functions only from sampled transitions (observed sets of state, action, reward, next-state) which only allows reduction of the sampled Bellman error at that state-action pair. This allows us to improve policies in domains where it is easier to build a simulator than specify the model of the environment and rewards e.g. board games, physics simulators for robots. Second, dynamic programming makes a “breadth-first” (accross all states and actions) update of the solution at each sweep, while reinforcement learning makes a “incremental” (only for the current state and action) update; this greatly reduces the computational burden and enhances scalability.
Reduction of average Bellman errors gives dynamic programming a geometric rate of convergence to the optimal solution, while the incremental updates and reduction of sampled Bellman errors, when combined with “sufficient exploration” of the state-action space, gives reinforcement learning only sublinear convergence guarantees. This is however, quite sufficient in practise, the scalability attained by only sampling transitions and making incremental updates more than makes up for the slower rates of convergence (and brittleness). These approximation methods sacrifice theoretical guarantees. RL algorithms lack the convergence assurances of exact dynamic programming. They exhibit sensitivity to hyperparameters and initialization. They can converge to suboptimal policies without diagnostic indication. This survey presents reinforcement learning as a computationally flexible framework while acknowledging its methodological limitations.
@@ 2 Two Cultures of Sequential Decision-Making
The two fields also differ in their treatment of human agency. To an economist, the human (or firm) is the optimizer, a rational or boundedly rational agent solving her own dynamic program, whose revealed choices are the data to be explained. To an RL researcher, the algorithm is the optimizer and the human is part of the environment. A customer’s purchasing decision is not modeled as utility maximization but as a stochastic reward signal from which the algorithm learns. Even the word “learning” itself carries different meaning across the divide. In economic theory, learning means Bayesian updating of beliefs about unknown parameters. In reinforcement learning, learning means executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator.
@@ 2.1 Core Distinctions
Aside from data access is the question of what the algorithm does with that data. In RL, the word model refers strictly to the environment’s dynamics, the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$. A model-based RL algorithm explicitly constructs or is given a mathematical representation of $P$ and $r$, then computes a policy by planning through that representation (for example, using a simulator or the known rules of a game, as in AlphaZero). A model-free algorithm computes the value function or policy directly from experienced transitions without ever building an explicit representation of the transition probabilities. Crucially, “model-free” need not mean the algorithm lacks complete access to any model of the environment. A model-free algorithm could interact with a simulator that internally implements a complete computerized model of the environment; the distinction is that the algorithm never extracts or plans through the transition probabilities, treating the simulator as a black box that merely returns sample transitions. In some sense, usually when we use “model-free” learning we are saying that while we do have access to some model, for some reason we do not or simply can not use it to find an optimal policy directly and instead find it easier to “sample transitions and rewards” from it. So an agent may have access to a complete simulation model of the world (e.g., a chess engine), but if its algorithm learns purely from sampled trajectories rather than using the known rules of the game to plan, it is considered model-free learning. However, it is possible that a “model-free” algorithm could also be implimented “in-field” where there is genuinely no access to a “model” of the environment but only direct access to the environment itself (for reasons discussed later, this is rarely done).
@@
One might believe that an analog to “model-free” RL algorithms might be reduced-form or a statistical model that only requires a few assumptions (exogeneity, etc.) to make inference on parameters of interest. But this is a false analogy, as regardless of “model-free” or “model-based” we are talking about learning optimal policies within a MDP (this itself demands a certain structure on the problem including the Markovian assumption that the state contains all relevant information). One could believe that “model-based” algorithms are analogous to structural estimation, which involves the specification of preferences, technology, beliefs, and an equilibrium concept, all of which give rise to specific transition dynamics and reward functions. This is also a false analogy. In RL terminology, for an algorithm to be “model-based”, the conditions very precise, all we need is a given representation of $P$ and $r$ (regardless of whether you do or don’t make any structural economic assumptions about human or firm behavior or their interactions). All this to say, in RL, the model generally only refers to transition matrices $P$ and rewards $r$ within an MDP; wheras in economics it refers to a set of agents, preferences, exclusion restrictions, and equilibrium concepts. Therefore, it is possible to do structural estimation of an “economic model” with a lot of structure but by using a “model-free” RL algorithm as a computational tool to find the equilibrium, as in Section 8.333The closer analog to structural estimation in the RL taxonomy is inverse reinforcement learning (IRL), which takes observed behavior as data and recovers the reward function that rationalizes it. This is the subject of the companion survey (Rust and Rawat, 2026).
@@
Every RL system passes through two phases. In the training phase, the agent interacts with an environment (simulated or real) and updates its parameters. In the execution phase (also called the deployment phase), the policy is frozen and used to make decisions without further updates. This distinction is critical for interpreting “online.” Online training in RL almost always takes place inside a simulator (and not in the “real world”). AlphaGo Zero trained online through millions of self-play games; when it faced Lee Sedol, its weights were frozen and it was purely executing its trained policy. Some deployed systems in Section 7 followed this pattern cleanly; DiDi’s dispatch system trained a value function from historical trip data, then deployed with fixed weights. Others blur the boundary. The hotel revenue management system in Section 7.2 updated Q-values from realized returns after each completed episode during live operation, making it an in-field learner rather than a frozen executor.444“In-field” is not standard RL vocabulary. We introduce it here to distinguish live-market online learning, where exploration has real economic cost, from the far more common case of online learning inside a simulator. Bandit algorithms (Section 10) also learn in-field by design, updating demand estimates from real customer responses during deployment. The term in-field is useful here; it denotes online learning where the environment is a live market with real economic consequences, as opposed to online learning inside a simulator where exploration is free, cheap and safe. While industry practitioners use terms like ’operational phase’, ’live’, and ’inference-time’ interchangeably to describe any post-training system, this survey reserves the term in-field to specifically denote online learning in a live market where exploration carries real economic consequences.
@@ 2.2 Overlapping Terminology
In RL the environment is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^5] A source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.
@@
Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988).
@@
Learning carries at least four distinct meanings in this survey. In economic theory (Bayesian learning, adaptive expectations), learning refers to agents forming and refining beliefs about unknown parameters of their environment. In supervised machine learning, learning means statistical estimation, fitting the weights of a parameterized model to minimize a loss function over data. In reinforcement learning, “learning” is primarily computation. When an RL agent “learns” a Q-function, it is executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator. We say the algorithm is “learning” because it improves its policy iteratively through simulated or real experience, but mathematically it is solving a fixed-point problem. In many applications throughout this survey, particularly Section 8, the RL algorithm is simply a numerical method for solving the Bellman equation; no actual human-like learning from experience is taking place. Finally, while RL draws inspiration from animal psychology (Section 3.1), it is a drastic simplification of biological learning. Tabula rasa RL algorithms require millions of iterations of trial and error to discover policies that animals acquire rapidly. Real-world animal learning relies on innate priors, basic physical knowledge, and parental nurturing and should be distinct from RL-style “learning”.
Function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation, the approximation of an unknown function by projection onto a finite-dimensional basis.
@@
A concrete example clarifies how the two culture see a common problem. The RL literature frames some recommender systems as contextual bandits, where user covariates are the context, the recommendation is the arm, and a click or rating is the reward. An economist might instead view movie recommendation as a two-sided learning problem in which the platform learns user preferences while users simultaneously explore the catalog and update their own tastes. The bandit formulation absorbs the user’s utility maximization into the environment’s reward signal and treats user arrivals as exogenous. This is a modeling choice, not a fact about the world. Also, the object called a “bandit” in this formulation, a one-step decision under exogenous context, is a slightly different mathematical object than the Bayesian sequential allocation problem of Gittins (1979), even though both carry the same name and are related.  [^7] Lattimore (2016) proves that the Gittins index with a flat Gaussian prior achieves finite-time regret of the same order as UCB, and that the index decomposes as posterior mean plus an exploration bonus that shrinks toward zero near the horizon, structurally resembling but differing from the UCB confidence bound. As apparant, the RL view often abstracts away from a lot of complexity in the world (human learning, human interaction) to squash the problem into the straightforward language of optimal control ($pi$, $V$, $P,r$), while the economics literature prefers to allow for much more complex models of human-firm interaction.

### 23 Mar 2026 v4

@@ A Survey of Reinforcement Learning For Economics
I am grateful to John Rust for encouraging me to undertake this survey. I thank Simon LeBastard for sharing his RL notes with me.
@@ 1 Introduction
This survey (re)introduces reinforcement learning to economists. I review the theoretical connections between dynamic programming and reinforcement learning, demonstrating how value iteration, Q-learning, and policy gradient methods are common solution methods to the same class of optimization problems. I then examine applications across several domains, including control problems, pricing and inventory management; structural economic models with high-dimensional state spaces; strategic games in which multi-agent algorithms compute equilibria under imperfect information; bandit problems in which economic structure yields tighter regret bounds; and preference learning. The exposition combines formal theory, practical applications and computational illustration.
Both dynamic programming and reinforcement learning solve the Bellman equation; they differ in the information requirements and the way in which the solution is refined. First, dynamic programming requires knowledge of the transition in the environment and the reward function which allows the reduction of the average Bellman error, reinforcement learning estimates value functions only from sampled transitions (observed sets of state, action, reward, next-state) which only allows reduction of the sampled Bellman error at that state-action pair. This allows us to improve policies in domains where it is easier to build a simulator than specify the model of the environment and rewards e.g. board games, physics simulators for robots. Second, dynamic programming makes a “breadth-first” (across all states and actions) update of the solution at each sweep, while reinforcement learning makes a “incremental” (only for the current state and action) update; this greatly reduces the computational burden and enhances scalability.
Reduction of average Bellman errors gives dynamic programming a geometric rate of convergence to the optimal solution, while the incremental updates and reduction of sampled Bellman errors, when combined with “sufficient exploration” of the state-action space, gives reinforcement learning only sublinear convergence guarantees. This is however, quite sufficient in practice, the scalability attained by only sampling transitions and making incremental updates more than makes up for the slower rates of convergence (and brittleness). These approximation methods sacrifice theoretical guarantees. RL algorithms lack the convergence assurances of exact dynamic programming. They exhibit sensitivity to hyperparameters and initialization. They can converge to suboptimal policies without diagnostic indication. This survey presents reinforcement learning as a computationally flexible framework while acknowledging its methodological limitations.
@@ 2 Two Cultures of Sequential Decision-Making
The two fields also differ in their treatment of human agency and in the meaning of basic terms like “agent,” “learning,” “model,” and “policy,” as detailed in the subsections that follow.
@@ 2.1 Core Distinctions
Aside from data access is the question of what the algorithm does with that data. In RL, the word model refers strictly to the environment’s dynamics, the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$. A model-based RL algorithm explicitly constructs or is given a mathematical representation of $P$ and $r$, then computes a policy by planning through that representation (for example, using a simulator or the known rules of a game, as in AlphaZero). A model-free algorithm computes the value function or policy directly from experienced transitions without ever building an explicit representation of the transition probabilities. Crucially, “model-free” need not mean the algorithm lacks complete access to any model of the environment. A model-free algorithm could interact with a simulator that internally implements a complete computerized model of the environment; the distinction is that the algorithm never extracts or plans through the transition probabilities, treating the simulator as a black box that merely returns sample transitions. However, it is possible that a “model-free” algorithm could also be implemented “in-field” where there is genuinely no access to a “model” of the environment but only direct access to the environment itself (for reasons discussed later, this is rarely done).
@@
One might believe that an analog to “model-free” RL algorithms might be reduced-form or a statistical model that only requires a few assumptions (exogeneity, etc.) to make inference on parameters of interest. But this is a false analogy, as regardless of “model-free” or “model-based” we are talking about learning optimal policies within a MDP (this itself demands a certain structure on the problem including the Markovian assumption that the state contains all relevant information). One could believe that “model-based” algorithms are analogous to structural estimation, which involves the specification of preferences, technology, beliefs, and an equilibrium concept, all of which give rise to specific transition dynamics and reward functions. This is also a false analogy. In RL terminology, for an algorithm to be “model-based”, the conditions are very precise, all we need is a given representation of $P$ and $r$ (regardless of whether you do or don’t make any structural economic assumptions about human or firm behavior or their interactions). All this to say, in RL, the model generally only refers to transition matrices $P$ and rewards $r$ within an MDP; whereas in economics it refers to a set of agents, preferences, exclusion restrictions, and equilibrium concepts. Therefore, it is possible to do structural estimation of an “economic model” with a lot of structure but by using a “model-free” RL algorithm as a computational tool to find the equilibrium, as in Section 8.
@@
Every RL system passes through two phases. In the training phase, the agent interacts with an environment (simulated or real) and updates its parameters. In the execution phase (also called the deployment phase), the policy is frozen and used to make decisions without further updates. This distinction is critical for interpreting “online.” Online training in RL almost always takes place inside a simulator (and not in the “real world”). AlphaGo Zero trained online through millions of self-play games; when it faced Lee Sedol, its weights were frozen and it was purely executing its trained policy. Some deployed systems in Section 7 followed this pattern cleanly; DiDi’s dispatch system trained a value function from historical trip data, then deployed with fixed weights. Others blur the boundary. The hotel revenue management system in Section 7.2 updated Q-values from realized returns after each completed episode during live operation, making it an in-field learner rather than a frozen executor.333“In-field” is not standard RL vocabulary. We introduce it here to distinguish live-market online learning, where exploration has real economic cost, from the far more common case of online learning inside a simulator. Bandit algorithms (Section 10) also learn in-field by design, updating demand estimates from real customer responses during deployment.
@@ 2.2 Overlapping Terminology
The word agent refers to different entities. In economics, an agent is always a human decision-maker (a consumer, worker, or firm) or a social planner whose choices are the object of study. In RL, “the agent” is the learning algorithm itself, or more precisely an algorithm deployed on behalf of a human decision-maker, and the human, if present at all, is part of the environment providing reward signals.
In RL the environment is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^4] A source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.. RL speaks of rewards where economists speak of utility. The mapping is not exact: a reward $r(s,a)$ is a known, externally specified function that the algorithm maximizes, whereas utility $u(x,d)$ in economics is a primitive of preferences that the analyst must recover or estimate from observed choices.
@@
Learning carries many distinct meanings in the two fields. In economic theory (Bayesian learning, adaptive expectations), learning refers to agents forming and refining beliefs about unknown parameters of their environment. In supervised machine learning, learning means statistical estimation, fitting the weights of a parameterized model to minimize a loss function over data. In reinforcement learning, “learning” is primarily computation. When an RL agent “learns” a Q-function, it is executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator. We say the algorithm is “learning” because it improves its policy iteratively through simulated or real experience, but mathematically it is solving a fixed-point problem. In many applications throughout this survey, particularly Section 8, the RL algorithm is simply a numerical method for solving the Bellman equation; no actual human-like learning from experience is taking place. Finally, while RL draws inspiration from animal psychology (Section 3.1), it is a drastic simplification of biological learning. Tabula rasa RL algorithms require millions of iterations of trial and error to discover policies that animals acquire rapidly. Real-world animal learning relies on innate priors, basic physical knowledge, and parental nurturing and should be distinct from RL-style “learning”.
Macroeconomic adaptive learning has used ideas formally analogous to reinforcement learning for decades, though under different names and with different motivations. In the adaptive learning literature initiated by Marcet and Sargent (1989), boundedly rational agents update their beliefs about equilibrium parameters using recursive least-squares and other Robbins-Monro-type stochastic approximation algorithms. The convergence criterion in that literature, E-stability, evaluates whether the ordinary differential equation (ODE) associated with the mapping from the perceived to the actual law of motion is locally asymptotically stable at the rational expectations equilibrium. This fixed-point stability requirement is analogous to the contraction conditions governing the convergence of temporal-difference and Q-learning algorithms in RL. Borkar and Meyn (2000) make this mathematical connection explicit: they prove the convergence of Q-learning and actor-critic methods using the ODE method, explicitly citing Sargent (1993) as a parallel application of the exact same framework to boundedly rational economic agents.
Active learning appears in both fields but means different things. In economics, active learning denotes Bayesian experimentation where an agent endogenously chooses what information to acquire, trading off the cost of exploration against the option value of better future decisions. In machine learning, active learning is a supervised learning protocol in which the algorithm selects which unlabeled examples to query an oracle for labels, minimizing annotation cost rather than maximizing cumulative reward. The term exploration is ubiquitous in RL but rarely used in economics. In RL, exploration is a modular subcomponent of a larger algorithm, a procedure for choosing informative actions that may be quite crude (such as $\varepsilon$-greedy random action selection) or more principled (UCB, Thompson sampling). The closest economic analog is optimal experimentation in the Bayesian bandit tradition (Rothschild 1974), where the value of information is derived endogenously from the agent’s dynamic program, not bolted on as a separate heuristic. Economists also use Bayesian learning to describe the same phenomenon, but always within a fully specified model of beliefs and preferences; in RL, exploration can be a purely algorithmic device with no decision-theoretic foundation.
Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988). Function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation, the approximation of an unknown function by projection onto a finite-dimensional basis. Calibration carries unrelated meanings. In economics, calibration means choosing model parameters to match a set of empirical moments or stylized facts (Kydland and Prescott 1982). In machine learning, calibration refers to probability calibration, ensuring that predicted probabilities match observed frequencies, a statistical property of a classifier’s outputs.
@@
A concrete example clarifies how the two culture see a common problem. The RL literature frames some recommender systems as contextual bandits, where user covariates are the context, the recommendation is the arm, and a click or rating is the reward. An economist might instead view movie recommendation as a two-sided learning problem in which the platform learns user preferences while users simultaneously explore the catalog and update their own tastes. The bandit formulation absorbs the user’s utility maximization into the environment’s reward signal and treats user arrivals as exogenous. This is a modeling choice, not a fact about the world. Also, the object called a “bandit” in this formulation, a one-step decision under exogenous context, is a slightly different mathematical object than the Bayesian sequential allocation problem of Gittins (1979), even though both carry the same name and are related.  [^6] Lattimore (2016) proves that the Gittins index with a flat Gaussian prior achieves finite-time regret of the same order as UCB, and that the index decomposes as posterior mean plus an exploration bonus that shrinks toward zero near the horizon, structurally resembling but differing from the UCB confidence bound. As apparent, the RL view often abstracts away from a lot of complexity in the world (human learning, human interaction) to squash the problem into the straightforward language of optimal control ($\pi$, $V$, $P,r$), while the economics literature prefers to allow for much more complex models of human-firm interaction.
The word policy is itself overloaded, and the divergence runs deeper than terminology. In economics, a “policy” is a rule set by a government, central bank, or regulator, a tax schedule, a subsidy, an interest rate rule, a licensing requirement. These rules are part of the environment in which private agents (consumers, firms) optimize. “Policy evaluation” in this tradition means changing some aspect of the environment and asking how agents would respond: if the earned income tax credit were expanded, how would labor supply shift? The standard workflow proceeds in two steps: first estimate or calibrate the structural model (preferences, technology, market interactions), then simulate counterfactuals under the alternative rule. In RL, “policy” means the agent’s own decision rule $\pi(a|s)$, and “policy evaluation” means computing $V^{\pi}(s)$, the expected return from following that rule in a fixed environment. RL typically assumes access to a high-fidelity simulator, a physics engine, a game, a digital twin, whose rules do not change; the interesting question is how the agent should behave within those rules, not what happens if the rules change. When the environment does shift, RL treats it as a nuisance (sim-to-real transfer, domain adaptation) rather than the object of study; in offline RL (Section 11), distributional shift between the training data and the deployed environment becomes a central concern, bringing RL closer to the economist’s counterfactual reasoning. [^7] A notable exception is Tomasev et al. (2020), where AlphaZero was used to evaluate alternative chess rule sets proposed by Kramnik, asking how optimal play changes under modified game rules. This is precisely an environment counterfactual of the kind economists routinely study.
Identification carries fundamentally different weight across the two fields. In econometrics, a parameter $\theta$ is identified if it is uniquely pinned down by the combination of the data-generating process and the model’s maintained assumptions; formally, no two distinct parameter values $\theta\neq\tilde{\theta}$ can generate the same distribution of observable data (Lewbel, 2019). Identification is a logical property of the model, not a statistical one; if identification fails, no amount of additional data or more sophisticated estimation will recover the parameter, because multiple parameter values are observationally equivalent. In RL, there is little general identification discourse. The closest analogs are narrow and domain-specific. In inverse reinforcement learning, reward identifiability asks whether the reward function can be uniquely recovered from observed optimal behavior; Kim et al. (2021) formalize this for MaxEnt MDP models, showing that for deterministic dynamics, strong identifiability requires the domain graph to be coverable and aperiodic. [^8] Kim et al. (2021) explicitly note that the econometrics literature on identifying dynamic discrete choice models (Rust, 1994) addresses “an equivalent problem” to reward identifiability in IRL, providing a direct bridge between the two fields. In model-based RL, system identification refers to learning the transition dynamics $\hat{T}$ well enough for the resulting policy to perform well, a usage inherited from control theory rather than econometrics (Ross and Bagnell, 2012). RL researchers focus on out-of-sample performance of the learned controller, so whether parameters are identified in the econometric sense matters less; what matters is that the policy generalizes. [^9] The one RL subfield where identification is central, inverse reinforcement learning, is the subject of the companion survey (Rust and Rawat, 2026).
The term regret itself diverges across fields. In economics, regret is either an emotion that modifies preferences under uncertainty (Loomes and Sugden 1982) or a static minimax decision criterion for choosing among actions when probabilities are unknown (Savage 1951, Manski 2004). In RL and computer science, regret is a dynamic performance metric, the cumulative gap between the agent’s returns and those of the best fixed policy in hindsight, and sublinear growth of this quantity is the standard benchmark for online learning algorithms. In economics, efficiency concerns welfare: Pareto efficiency asks whether any reallocation could improve one agent’s utility without harming another, and allocative efficiency asks whether goods flow to their highest-valued uses. In RL, efficiency concerns resources: sample efficiency measures how many environment interactions an algorithm needs to find a good policy, and computational efficiency measures the operations required per timestep.
In economics, the discount factor is a structural parameter encoding time preference, an agent’s intrinsic willingness to trade present for future consumption. Koopmans (1960) derived it axiomatically from preference postulates, principally stationarity and impatience, showing that these behavioral axioms uniquely imply an exponential discounted utility representation with a single discount factor strictly between zero and one (Bleichrodt et al., 2008). Economists treat this parameter as a deep primitive of preferences, to be estimated from data on intertemporal choices, and deviations from constant discounting (such as the quasi-hyperbolic present bias of Laibson 1997) are studied as substantive behavioral phenomena. In RL, the discount factor has historically served a more pragmatic role, ensuring that infinite-horizon returns remain finite and that the Bellman operator is a contraction, thereby guaranteeing convergence of dynamic programming algorithms. It is typically treated as a hyperparameter to be tuned for computational performance rather than a structural claim about the agent’s preferences. Pitis (2019) bridges this gap, axiomatically deriving discounting in RL from rationality postulates and showing that the fixed discount factor is best understood as an optimizing representation rather than a literal description of time preference.